# Gradient-free stochastic optimization for additive models

### v1

@@ 2 Problem setup
In order to minimize $f$, we apply a version of projected gradient descent with estimated gradient defined as follows. Let $\{\eta_{t}\}_{t=1}^{T}$ be a sequence of positive numbers and let $\{\boldsymbol{g}_{t}\}_{t=1}^{T}$ be a sequence of random vectors. Consider any fixed $\boldsymbol{x}_{1}\in\mathbb{R}^{d}$ and let the vectors $\boldsymbol{x}_{t}$ for $t=2,\dots,T$ be defined by the recursion
@@
We set $\boldsymbol{g}_{t}=(g_{t,1},\dots,g_{t,d})$.
@@ 3 Statement of the results
In this section, we provide upper bounds on the optimization error of the algorithm defined in Section 2. First, we assume that function $f$ represented by the additive model satisfies the $\alpha$-PL condition. Note that imposing this condition on the sum $f$ does not imply that all the components $f_{j}$ are PL functions. When dealing with PL functions we consider the problem of unconstrained minimization and we introduce the notation
@@
Let $\alpha>0$, $\beta\geq 2$ and $\bar{L},L>0$. For $j\in[d]$, let $f_{j}:\mathbb{R}\to\mathbb{R}$ be such that $f_{j}\in\mathcal{F}_{2}^{\prime}(\bar{L})\cap\mathcal{F}_{\beta}(L)$. Assume that $f:\mathbb{R}^{d}\to\mathbb{R}$ has the form $f(\boldsymbol{x})=\sum_{j=1}^{d}f_{j}(x_{j})$, and that $f$ is an $\alpha$-PL function. Let Assumption 1 hold, and let $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ be the updates of algorithm (2) with $\Theta=\mathbb{R}^{d}$, where the gradient estimators $\boldsymbol{g}_{t}$ are defined by (4), and
@@
Let $\alpha>0$, $\beta\geq 2$ and $\bar{L},L>0$. For $j\in[d]$, let $f_{j}:\mathbb{R}\to\mathbb{R}$ be such that $f_{j}\in\mathcal{F}_{2}^{\prime}(\bar{L})\cap\mathcal{F}_{\beta}(L)$. Assume that $f:\mathbb{R}^{d}\to\mathbb{R}$ has the form $f(\boldsymbol{x})=\sum_{j=1}^{d}f_{j}(x_{j})$, and that $f$ is an $\alpha$-strongly convex function such that $\max_{\boldsymbol{x}\in\Theta}\left\lVert\nabla f(\boldsymbol{x})\right\rVert\leq G$. Let Assumption 1 hold, and let $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ be the updates of algorithm (2), where the gradient estimators $\boldsymbol{g}_{t}$ are defined by (4) with
@@
Consider the weighted updates
@@ 4 Proofs
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be a differentiable function such that $\left\lVert\nabla f(\boldsymbol{x})-\nabla f(\boldsymbol{x}^{\prime})\right\rVert\leq\bar{L}\left\lVert\boldsymbol{x}-\boldsymbol{x}^{\prime}\right\rVert$ for all $\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathbb{R}^{d}$, where $\bar{L}>0$. Let Assumption 1 hold and let $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ be the updates of algorithm (2) with $\Theta=\mathbb{R}^{d}$. Then, for all $t\in[T]$ we have
@@
$|\mathbf{E}\left[g_{t,j}|x_{t,j}\right]-f^{\prime}_{j}(x_{t,j})|\leq L\mathbf{E}\left[|r_{t,j}|^{\beta}K(r_{t,j})\right]h_{t}^{\beta-1}=\kappa_{\beta}Lh_{t}^{\beta-1},$
and thus the result of the lemma. ∎
@@
For $j\in[d]$, let $f_{j}:\mathbb{R}\to\mathbb{R}$ be such that $f_{j}\in\mathcal{F}_{2}^{\prime}(\bar{L})\cap\mathcal{F}_{\beta}(L)$, where $\beta\geq 2$ and $\bar{L},L>0$. Assume that $f:\mathbb{R}^{d}\to\mathbb{R}$ satisfies the additive model $f(\boldsymbol{x})=\sum_{j=1}^{d}f_{j}(x_{j})$. Let Assumption 1 hold. Assume that $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ are updates of algorithm (2) with $\Theta=\mathbb{R}^{d}$ and with the gradient estimators defined by (4). Then, for all $t\in[T]$ we have
@@
For $\alpha>0$ assume that $f$ is an $\alpha$-strongly convex function. Assume that $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ are the updates of algorithm (2). Then, for all $t\in[T]$ and $\boldsymbol{x}\in\Theta$, we have

### v2

@@ 2 Problem setup
Input: Constraint set $\Theta$, kernel function $K:[-1,1]\to\mathbb{R}$, step size $\eta_{t}>0$ and perturbation parameter $h_{t}>0$, for $t=1,\dots,T$
Initialization: Generate vectors $\boldsymbol{r}_{t}=(r_{t,1},\dots,r_{t,d})\in\mathbb{R}^{d}$ for $t=1,\dots,T$, where the components $r_{t,i}$ are independent and distributed uniformly from the interval $[-1,1]$, and choose $\boldsymbol{x}_{1}\in\Theta$
for $t=1,\dots,T$ do
       Observe $y_{t}=f(\boldsymbol{x}_{t}+h_{t}\boldsymbol{r}_{t})+\xi_{t}$ and $y^{\prime}=f(\boldsymbol{x}_{t}-h_{t}\boldsymbol{r}_{t})+\xi_{t}^{\prime}$
        // query
       for $j=1,\dots,d$ do
             $g_{t,j}=\frac{d}{2h_{t}}(y_{t}-y_{t}^{\prime})K(r_{t,j})$
       end for
      Let $\boldsymbol{g}_{t}=(g_{t,1},\dots,g_{t,d})$
        // gradient estimator
       $\boldsymbol{x}_{t+1}=\text{\rm Proj}_{\Theta}(\boldsymbol{x}_{t}-\eta_{t}\boldsymbol{g}_{t})$
        // update
end for
Algorithm 1 Zero-Order Stochastic Projected Gradient
In order to minimize $f$, we apply a version of projected gradient descent with estimated gradient presented in Algorithm 1. Let $\{\eta_{t}\}_{t=1}^{T}$ be a sequence of positive numbers and let $\{\boldsymbol{g}_{t}\}_{t=1}^{T}$ be a sequence of random vectors. Consider any fixed $\boldsymbol{x}_{1}\in\mathbb{R}^{d}$ and let the vectors $\boldsymbol{x}_{t}$ for $t=2,\dots,T$ be defined by the recursion
@@
We consider the gradient estimator $\boldsymbol{g}_{t}=(g_{t,1},\dots,g_{t,d})$. Note that other choices of gradient estimator can lead to similar results as those that we obtain below, namely estimators based on finite difference approximations taking into account higher order smoothness. In contrast to (4), such higher order finite difference schemes have a complicated form and require many queries per step of the algorithm.
@@ 3 Statement of the results
In this section, we provide upper bounds on the optimization error of the algorithm defined in Section 2. First, we assume that function $f$ represented by the additive model satisfies the $\alpha$-PL condition. Note that imposing this condition on the sum $f$ implies that all the components $f_{j}$ are PL functions. When dealing with PL functions we consider the problem of unconstrained minimization and we introduce the notation
@@
Let $\alpha>0$, $\beta\geq 2$ and $\bar{L},L>0$. For $j\in[d]$, let $f_{j}:\mathbb{R}\to\mathbb{R}$ be such that $f_{j}\in\mathcal{F}_{2}^{\prime}(\bar{L})\cap\mathcal{F}_{\beta}(L)$. Assume that $f:\mathbb{R}^{d}\to\mathbb{R}$ has the form $f(\boldsymbol{x})=\sum_{j=1}^{d}f_{j}(x_{j})$, and that $f$ is an $\alpha$-PL function. Let Assumption 1 hold, and let $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ be the updates of Algorithm 1 with $\Theta=\mathbb{R}^{d}$, and
@@
Let $\alpha>0$, $\beta\geq 2$ and $\bar{L},L>0$. For $j\in[d]$, let $f_{j}:\mathbb{R}\to\mathbb{R}$ be such that $f_{j}\in\mathcal{F}_{2}^{\prime}(\bar{L})\cap\mathcal{F}_{\beta}(L)$. Assume that $f:\mathbb{R}^{d}\to\mathbb{R}$ has the form $f(\boldsymbol{x})=\sum_{j=1}^{d}f_{j}(x_{j})$, and that $f$ is an $\alpha$-strongly convex function on a compact convex subset $\Theta$ of $\mathbb{R}^{d}$ such that $\max_{\boldsymbol{x}\in\Theta}\left\lVert\nabla f(\boldsymbol{x})\right\rVert\leq G$. Let Assumption 1 hold, and let $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ be the updates of Algorithm 1 with
@@
Consider the weighted estimator
@@
Remark 1.
In view of strong convexity, Theorem 2 and Corollary 2 immediately imply the corresponding bounds on the estimation error $\mathbf{E}[\left\lVert\bar{\boldsymbol{x}}_{T}-\boldsymbol{x}^{*}\right\rVert^{2}]$, where $\boldsymbol{x}^{*}$ is the minimizer of $f$ on $\Theta$. Thus, under the assumptions of Corollary 2 we have
$\mathbf{E}[\left\lVert\bar{\boldsymbol{x}}_{T}-\boldsymbol{x}^{*}\right\rVert^{2}]\leq 2\texttt{A}\frac{d}{\alpha^{2}}T^{-\frac{\beta-1}{\beta}},$
where $\texttt{A}>0$ is the constant from Corollary 2.
@@
Remark 2.
Since the strong convexity and the PL property hold for each additive component of $f$, another possible approach would be to run the procedure component-wise (minimize separately each component $f_{j}$). However, it leads to a worse result. Indeed, in this case we need to make at each step $2d$ queries in parallel (two queries for each component) and thus can make only $\sim T/d$ steps if the total budget of queries is $T$. At the end, applying Theorems 1 or 2 in the one-dimensional case, for each component we obtain the error in $T$ and $d$ of the order $(T/d)^{-(\beta-1)/\beta}$. This rate cannot be improved as follows from the one-dimensional instance of Theorem 3. Summing up over the $d$ components, the overall error will be of the order $d(T/d)^{-(\beta-1)/\beta}=d^{2-1/\beta}T^{-(\beta-1)/\beta}$, that is, the error will depend on $d$ in a sub-optimal way.
@@ 4 Proofs
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be a differentiable function such that $\left\lVert\nabla f(\boldsymbol{x})-\nabla f(\boldsymbol{x}^{\prime})\right\rVert\leq\bar{L}\left\lVert\boldsymbol{x}-\boldsymbol{x}^{\prime}\right\rVert$ for all $\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathbb{R}^{d}$, where $\bar{L}>0$. Let Assumption 1 hold and let $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ be the updates of Algorithm 1 with $\Theta=\mathbb{R}^{d}$. Then, for all $t\in[T]$ we have
@@
$|\mathbf{E}\left[g_{t,j}|{\boldsymbol{x}_{t}}\right]-f^{\prime}_{j}(x_{t,j})|\leq L\mathbf{E}\left[|r_{t,j}|^{\beta}K(r_{t,j})\right]h_{t}^{\beta-1}=\kappa_{\beta}Lh_{t}^{\beta-1}.$
The result of the lemma follows from this inequality and the fact that
$\left\lVert\mathbf{E}\left[\boldsymbol{g}_{t}|\boldsymbol{x}_{t}\right]-\nabla f(\boldsymbol{x}_{t})\right\rVert\leq\sqrt{d}\max_{j=1,\dots,d}|\mathbf{E}\left[g_{t,j}|\boldsymbol{x}_{t}\right]-f^{\prime}_{j}(x_{t,j})|.$
@@
For $j\in[d]$, let $f_{j}:\mathbb{R}\to\mathbb{R}$ be such that $f_{j}\in\mathcal{F}_{2}^{\prime}(\bar{L})\cap\mathcal{F}_{\beta}(L)$, where $\beta\geq 2$ and $\bar{L},L>0$. Assume that $f:\mathbb{R}^{d}\to\mathbb{R}$ satisfies the additive model $f(\boldsymbol{x})=\sum_{j=1}^{d}f_{j}(x_{j})$. Let Assumption 1 hold. Assume that $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ are updates of Algorithm 1 with $\Theta=\mathbb{R}^{d}$ and with the gradient estimators defined by (4). Then, for all $t\in[T]$ we have
@@
For $\alpha>0$ assume that $f$ is an $\alpha$-strongly convex function. Assume that $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ are the updates of Algorithm 1. Then, for all $t\in[T]$ and $\boldsymbol{x}\in\Theta$, we have